OpenStudy (anonymous):

Which proportion has a solution of 8?

OpenStudy (texaschic101):

need the answer choices ?

OpenStudy (anonymous):

thirty over twelve equals n over forty-five twelve over thirty equals n over forty-five twelve over n equals forty-five over thirty twelve over thirty equals forty-five over n

OpenStudy (texaschic101):

let me show you how to solve a proportion... 30/12 = n/45 cross multiply (12)(n) = (45)(30) 12n = 1350 now divide by 12 to get n by itself 12n/12 = 1350/12 n = 112.5 obviously this is not the correct answer because n does not equal 8. you want to ry the next answer choice ?

OpenStudy (anonymous):

I need the answer in 5 mins

OpenStudy (texaschic101):

12/30 = n/45 30n = 540 n = 18 (no) 12/n = 45/30 45n = 360 n = 8 (BINGO, it is this one)

OpenStudy (anonymous):

thank you sooo much

OpenStudy (texaschic101):

no problem...do you understand how I did these ?

OpenStudy (anonymous):

sorta

OpenStudy (texaschic101):

cross multiply, then solve for the variable

OpenStudy (anonymous):

oh ok thank you is there some way you can like help me more often

OpenStudy (texaschic101):

when I am on, you can just tag me....as long as it is not geometry I am not good at that

OpenStudy (anonymous):

ok thanks

OpenStudy (texaschic101):

no problem :)